OpenStudy (anonymous):

4. One of the offices at the company was being repainted. Riley’s father challenged him to figure out how many quarts of paint would be needed. The office is 16 ft long, 12 ft wide, and the height of each wall is 10 ft. The office floor, ceiling, and door would not be painted and there are no windows in this office. The door opening measures 3 ft by 8 ft. One quart of paint covers 100 ft2. How many quarts of paint would be needed to paint the room, if only full quarts can be purchased? Show your work.

OpenStudy (annarf):

first, find out the total number of feet that need to be painted. Then, subtract unneeded area. Sqare feet of: wall 1 + 3 = 16 * 10 = 160 wall 2 + 4 = 12 * 10 = 120 total: 180 Subtract the door: door: 3 * 8 = 24 total after subtraction: 156 sq. ft. Now, to figure out how much paint is needed: 100 sq. ft = 1 quart. so 156 sq ft is about 1.5 quarts. However, we can only get full quarts, so we round up, to get 2 quarts of paint to cover the 156 sq ft of the office wall. Please tell me if that doesn't make sense! :)

OpenStudy (annarf):

:)